Judy is 12 years older than Joan. In 7 years, the sum of thier ages will be 42. How old is each now?
Answer by mananth(16946)

Judy is 12 years older than Joan. In 7 years, the sum of thier ages will be 42. How old is each now?
let joan's age be x
judy's age will be x+12
after 7 years their ages will be x+7 & (x+12) years
x+7 + ( x+12 )+7 = 42
x+7 + x+12 +7= 42
2x+26=42
2x=42-26
2x=12
x=6 which is joan"s age
Judy's age = 12+6 =18 years